OpenStudy (anonymous):

College Algebra: I'm doing 3X3 Linear Systems and I need help solving this problem. I tried a technique my professor taught me; however, it's not working this time. I need to use a system of equations to find the parabola of the form y=ax^2+bx+c that goes through the three given points: (0,4) (-2,0) (-3,1)

OpenStudy (nikato):

First, plug the x and y to that equation So you have 3 equations with 3 variables to solve(A,B, and C) And would like like to work on it using your professors technique?

OpenStudy (anonymous):

When I started plugging them in, I got c=4, 4a-2b+c=0, and 9a-3b+c=1. After that, everything I did didn't make sense

OpenStudy (nikato):

Ok. Everything's right so far. Can you show me what you did next?

OpenStudy (anonymous):

put c=4 ,then you are left with two equations in a and b solve them .

OpenStudy (anonymous):

Alright. Next, I plugged in the 4 where the C took place. So I had 4a-2b+4=0 and 9a-3b+4=1

OpenStudy (anonymous):

But of course, I moved it on the other side. So, it was 4a-2b=-4 and 9a-3b=-3

OpenStudy (nikato):

Good. Now you want to try to elimation one of the variable to solve for the other. For example ,multiplying both sides of the first equation by 3 and the other by 2 will get rid of b

OpenStudy (nikato):

And you'll be able to solve for what a is

OpenStudy (anonymous):

2a-b=-2 3a-b=-1 subtract and get the value of a and then b

OpenStudy (anonymous):

So, like this? \[-3(4a-2b)=(-4)-3\] \[-2(9a-3b)=(-3)-2\]

OpenStudy (anonymous):

Wait, what?

OpenStudy (nikato):

Exactly. And now multiply it out

OpenStudy (anonymous):

Alright, I got -12a+6b=12 and -18a+6b=6

OpenStudy (nikato):

Ok. Now subtract the two equations. Notice that the b's will cancel out. (6b-6b=0) And you should be able to solve for a

OpenStudy (anonymous):

a=1

OpenStudy (nikato):

Yea. And now you know a=1 and c=4. Just plug them into any of the equations to find b

OpenStudy (anonymous):

Oh cool, Alright. Let's see.

OpenStudy (anonymous):

b=4

OpenStudy (nikato):

Correct

OpenStudy (anonymous):

Ohhhh. So that's how the answer is y=x^2+4x+4

OpenStudy (anonymous):

Thank you so much Nikato. I truly do appreciate your help. :)

OpenStudy (nikato):

Yea. You're welcome. You were a good learner

OpenStudy (anonymous):

Thank you, I have a final exam coming up soon; therefore, I'm trying my best to pass. It's my best hope. :)

OpenStudy (nikato):

Good luck on your exam!

OpenStudy (anonymous):

Thank you!!!!